Solve for x:
4 + (16/3-x)=12

Answer by unlockmath(1688)

Hello,
For this 4 + (16/3-x)=12 we need to multiply by 3-x to get:
4(3-x) + 16=12(3-x)
Expand this out:
12-4x+16=36-12x
Add 12x and subtract 28:
8x=8
x=1
Make sense?
RJ
www.math-unlock.com